# Phases and phase transition in Grover’s algorithm with systematic noise

### v1

@@ III.2 Effective Hamiltonian of the bulk
We start by using our observation of clear linear $\delta$-dependence of the bulk spectrum to derive an effective Hamiltonian from first order degenerate perturbation theory. To achieve this goal, let us begin by writing the Grover operator with noise strength $\delta$ as
@@
where we have written each gate in the decomposition of the Grover operator as generalized rotations in Hilbert space (see Eq. (20)). Here $M=4L^{2}-6L+10$ is the total number of gates in the decomposition of the $L$-qubit Grover operator. Individual commuting gates are divided into separate terms in the product. Expanding the exponentials and keeping terms only up to first order in $\delta$, we obtain
@@
Thus, after discarding the two special states from $H_{\rm eff}$, its eigenvalues should match with the exact bulk quasienergies of $G(\delta)$. In Figure 4, we plot the noise-averaged difference between the exact and the effective energies:
@@
as a function of noise strength $\delta$. We see that deviations vanish linearly as $\delta\to 0$ , confirming the correctness of first-order perturbation theory. Indeed, we numerically see that first-order perturbation accurately captures the spectrum nearly up to the value of $\delta$ when the bulk spans the entire Floquet Brillouin zone from $-\pi$ to $\pi$.
Figure 4: Plot of noise averaged difference between the exact quasi-energy of $G(\delta)$ and eigenvalues of $\delta H_{\rm eff}$ for $L=8$ using Eq. (26). Error bands represents the standard error of mean for five noise realizations.
It is worth noting that our effective Hamiltonian is defined to be independent of the scaling parameter $\delta$, which allows us to analyze the statistical properties of $H_{\rm eff}$ for different noise strengths and therefore determine many important properties such as thermalization of the bulk states and scaling of $\delta_{c,\rm gap}$. Having shown the accuracy of $H_{\rm eff}$ in modeling our spectrum, we now turn to analyzing its properties with a focus on the bulk states. $H_{\rm eff}$ consists of random terms which start as 2-local terms before Heisenberg evolution causes operator spreading. Some remnant of the locality remains, as we argue numerically in the appendix (Section A). However, lacking any other structure, we would expect this Hamiltonian to exhibit generic, thermalizing behavior. We now show this by numerically demostrating its consistency with random matrix theory.
@@ III.2.1 Random matrix theory treatment of $H_{\mathrm{eff}}$
Figure 5: Plot of noise-averaged level spacing ratio (top) and Kullback-Leibler divergence (bottom) of $H_{\rm eff}$ for $L=10$. Error bars represents standard error of mean for $768$ noise realizations.
@@ III.3 Critical noise strength and its scaling
Having shown that $H_{\mathrm{eff}}$ obeys certain properties of RMT, we now use that to solve for other properties of the noisy algorithm, starting with the critical noise $\delta_{c,\rm gap}$ and its scaling with system size. We start by noting that the algorithm stops giving meaningful answers once the two special states enter into the bulk, i.e., once the gap closes. As evident from Figure 3, this happens near the point where the bulk energies span from $-\pi$ to $\pi$ due to weak dependence of the special state energies on $\delta$. We will study $\delta$-dependence of the special state energies in Section III.4. We define the noise value at this gap-closing point as the critical noise $\delta_{c,\rm gap}$. Since the quasi-energy of the noisy Grover operator is well captured by the effective Hamiltonian up to the point of merger of the special states, we will study the eigenvalues of the later to determine the critical noise and its scaling with system size.
@@
where we used that $\mathcal{E}\left(\epsilon_{j}\epsilon_{k}\right)=\frac{1}{3}\delta_{jk}$. Notice that $E_{0}\sim L$ comes from the overall gate count, $M\sim L^{2}$; by reducing scaling of the gate count, we would modify this system-size dependence. This scaling is corrobrated numerically in Figure 6.
@@
It is easy to see that irrespective of the fraction of eigenvalues we consider that span the entire Brillouin zone, the scaling of the critical noise does not change.
@@
for $1\leq m,n\leq N$, and where $\mathcal{E}_{\mathrm{GUE}}\left[\cdot\right]$ denotes the average over different realizations of the GUE and $\lambda$ sets the units of energy [31]. As a result of this discrepancy in scaling of elements, the distribution of the eigenvalues of $H_{\rm eff}$ deviates from the Wigner’s semicircle law, instead following a normal distribution
@@
Figure 6: $(a)$ Plot of the sum of squared diagonal and off-diagonal elements of $\tilde{H}_{\rm eff}$ as a function of system size in a log-log plot. Error bars indicate standard error of mean over $512$ noise realizations. The best fit functions $f(x)=aL^{2}$ (off-diagonal) and $g(L)=bL$ (diagonal) are shown with red and blue dashed lines respectively. $(b)$ (Left) $E_{0}$ as a function of system size $L$ calculated using Eq. (28) (red) by averaging over $512$ noise realizations, and is compared with the theoretical prediction given by (31). (Right) Distribution of bulk eigenvalues of $H_{\rm eff}$ for $L=10$, shown along with the normal distribution (Eq. (34)) (red) and Wigner’s semicircle (dashed).
Figure 7: Critical noise strength $\delta_{c,\rm gap}$ as a function of system size shown in a log-log plot. The exact critical noise is calculated at the point where the gap between the two special states and the bulk closes, and are averaged over $20$ noise realizations. The dashed line on the plot represents the best fit line for these exact noises. Predicted critical noises are averaged over $512$ noise realizations and are calculated when the two special states merges into $67\%,(-2E_{0},2E_{0})$, and $99.7\%,(-3E_{0},3E_{0})$ of the bulk eigenvalues spanning the Floquet Brillouin zone.
It is worth noting that the exact value of the gap closing between bulk and special states will not happen precisely at $\delta_{c,\rm gap}$ derived above, but rather a slightly smaller value. This is because the edges of the spectrum of $H_{\mathrm{eff}}$ do not scale identically as its variance. For example, conventional local models have extensive energy range but have energy width scaling as $\sqrt{L}$. However, the energy width $E_{0}\sim L$ is the most important parameter for determining ergodicity, as this is the energy scale on which a finite density of states will come in contact with the special states. The precise scaling of the range of bulk energies, which determines initial gap closing but not ergodicity breaking, will be left for future work.
@@ III.4 Analysis of the special states
Finally, we turn our attention to the two special states. In the absence of noise, the actions of the Grover operator on the states $|0\rangle$ and $|\bar{x}\rangle$ are
$G_{0}|0\rangle$ $=$ $\left(\frac{2}{N}-1\right)|0\rangle+\frac{2\sqrt{N-1}}{N}|\bar{x}\rangle$ (35)
$=$ $-\cos(2\theta)|0\rangle+\sin(2\theta)|\bar{x}\rangle,$ (36)
$G_{0}|\bar{x}\rangle$ $=$ $-\frac{2\sqrt{N-1}}{N}|0\rangle+\left(\frac{2}{N}-1\right)|\bar{x}\rangle$ (37)
$=$ $-\cos(2\theta)|\bar{x}\rangle-\sin(2\theta)|0\rangle.$ (38)
Then we can write the noiseless Grover operator in the $\{|0\rangle,|\bar{x}\rangle\}$ block as
@@
Now, adding back in the systematic noise, we can repeat our analysis using the effective Hamiltonian within first order perturbation theory and apply it to these special states. Restricting Eq.(25) to the special states manifold by writing it in $|\bar{0}\rangle$ and $|\bar{x}\rangle$ basis,
$H^{\rm spec}_{\rm eff}\equiv\begin{bmatrix}\langle 0|H_{\rm eff}|0\rangle&\langle 0|H_{\rm eff}|\bar{x}\rangle\\ \langle\bar{x}|H_{\rm eff}|0\rangle&\langle\bar{x}|H_{\rm eff}|\bar{x}\rangle\end{bmatrix},$ (40)
@@
The eigenvalues of $-2\theta\tau_{y}+\delta H^{\rm spec}_{\rm eff}-\pi$ capture the quasi-energies of two special states, as shown in Figure 8, once again confirming the effectiveness of first-order perturbation theory.
@@
Figure 9: (a) Variance of the coefficients $b_{x},b_{y}$ and $b_{z}$ as a function of system size shown along with their best fit curves. Data is averaged over $512$ noises. (b) Plot of average energy gap squared $\langle\Delta^{2}\rangle$ obtained using Eq. (43) as a function of system size $L$ for different noise strengths. Data is averaged over $512$ noise realizations.
If $H_{\mathrm{eff}}$ were an $N\times N$ GUE random matrix, then $H^{\rm spec}_{\rm eff}$ would be a $2\times 2$ GUE random matrix, for which $\mathcal{E}(b_{x,y,z}^{2})\sim 1/N$. However, as we have seen that non-GUE scaling of the diagonals exists in $H_{\mathrm{eff}}$, we need to more carefully examine whether GUE scaling holds within the special state manifold. As shown in Figure 9, indeed we do not have exponential scaling of these Pauli coefficients nor the gap. While the off-diagonal terms do decrease exponentially fast $\theta^{2},b^{2}_{x},b^{2}_{y}\sim 1/N$, the diagonal term instead increases as $b^{2}_{z}\sim L$. Putting these scalings into Eq. 43, we see that $\mathcal{E}\left[\Delta^{2}\right]\sim A/N+B\delta^{2}/N+CL\delta^{2}$ for some order-1 coefficients $A$, $B$, and $C$. Then, for $\delta\gg\delta_{c,\rm comp}\equiv\sqrt{A/(CLN)}\sim 1/\sqrt{L2^{L}}$, the linear $b_{z}$ term dominates over the $\theta^{2}$ term giving a power-law-large gap. While this increases the oscillation frequency within the special states, it comes at the cost of exponentially decreasing the oscillation amplitude. In the limit of $\delta\gg\delta_{c,\rm comp}$, the states $|0\rangle$ and $|\bar{x}\rangle$ become eigenstates of $G^{\rm spec}(\delta)$ and therefore their probabilities remain at
@@
at all times $T$, up to exponentially small oscillations. Since there is no longer an order-1 probability of finding the target state at any times, this implies that the model has lost computing power. Therefore, our special state analysis unfortunately implies that at this much smaller critical noise $\delta_{c,\rm comp}\sim L^{-1/2}2^{-L/2}$ we have a “computing transition,” because the algorithm loses computing power for $\delta\apprge\delta_{c,\rm comp}$. By contrast, the gap-closing critical noise $\delta_{c,\rm gap}\sim L^{-1}$ is much larger, but occurs well after computational power has been lost.
Given that the large coefficient $b_{z}$ in $H^{\rm spec}_{\rm eff}$ implies a special basis for the special state manifold, one might hope to recover an exponentially small gap by randomly rotating the basis, e.g., by choosing a different noise model. However, we now argue that this is not possible for any local or $k$-local noise model. The key point is that the relatively larger magnitude of the $b_{z}$ component compared to $b_{x}$ and $b_{y}$ comes from local distinguishability of the states $|0\rangle$ and $|\bar{x}\rangle$, meaning that generic local (or $k$-local) perturbations in $H_{\mathrm{eff}}$ will have a (random) diagonal component in this basis. By contrast, distinguishing the off-diagonal cat states $|0\rangle+e^{i\varphi}|\bar{x}\rangle$ requires global operators. This implies that $\left|\langle 0|H_{\rm eff}|\bar{x}\rangle\right|^{2}\ll\left|\langle 0|H_{\rm eff}|0\rangle\right|^{2},\left|\langle\bar{x}|H_{\rm eff}|\bar{x}\rangle\right|^{2}$, which produces $b_{z}^{2}\gg b_{x,y}^{2}$ for generic $k$-local error models.

### v2

@@ III.2 Effective Hamiltonian of the bulk
We start by using our observation of clear linear $\delta$-dependence of the bulk spectrum to derive an effective Hamiltonian from first-order degenerate perturbation theory. To achieve this goal, let us begin by writing the Grover operator with noise strength $\delta$ as
@@
where we have written each gate in the decomposition of the Grover operator as generalized rotations in Hilbert space [see Eq. (20)]. Here $M=4L^{2}-6L+10$ is the total number of gates in the decomposition of the $L$-qubit Grover operator. Individual commuting gates are divided into separate terms in the product. Expanding the exponentials and keeping terms only up to first order in $\delta$, we obtain:
@@
Thus, after discarding the two special states from $H_{\rm eff}$, its eigenvalues should match with the exact bulk quasienergies of $G(\delta)$. In Fig 4, we plot the noise-averaged difference between the exact and the effective energies:
@@
as a function of noise strength $\delta$. For small $\delta$, the $k-$th quasienergy can be written as $\phi_{\textrm{exact},k}=\delta E_{\textrm{eff},k}+\mathcal{O}(\delta^{2})$. From this, we can write $\phi_{\textrm{exact},k}/\delta-E_{\textrm{eff},k}=\mathcal{O}(\delta^{2})/\delta$. Then plotting $\phi_{\textrm{exact},k}/\delta-E_{\textrm{eff},k}$ as a function of $\delta$ will asymptote to zero as $\delta\to 0$ if the higher order terms $\delta^{3},\delta^{4}$ are small. In Fig 4 we plot the disorder averaged sum of all these $k-$quasienergies. We see that deviations vanish linearly as $\delta\to 0$, confirming the correctness of first-order perturbation theory. Indeed, we numerically see that first-order perturbation accurately captures the spectrum nearly up to the value of $\delta$ when the bulk spans the entire Floquet Brillouin zone from $-\pi$ to $\pi$.
Figure 4: Plot of noise averaged difference between the exact quasi-energy of $G(\delta)$ and eigenvalues of $\delta H_{\rm eff}$ for $L=8$ using Eq. (26). Error bands represent the standard error of mean for five noise realizations.
It is worth noting that our effective Hamiltonian is defined to be independent of the scaling parameter $\delta$, which allows us to analyze the statistical properties of $H_{\rm eff}$ for different noise strengths and therefore determine many important properties such as thermalization of the bulk states and scaling of $\delta_{c,\rm gap}$. Having shown the accuracy of $H_{\rm eff}$ in modeling our spectrum, we now turn to analyzing its properties with a focus on the bulk states. $H_{\rm eff}$ consists of random terms which start as 2-local terms before Heisenberg evolution causes operator spreading. Some remnant of the locality remains, as we argue numerically in Appendix A. However, lacking any other structure, we would expect this Hamiltonian to exhibit generic, thermalizing behavior. An exact analytical treatment of this Hamiltonian is not feasible, therefore we resort to analyzing its properties statistically using random matrix theory, which is a natural and powerful framework to study disordered many-body systems. One way to test if the Hamiltonian belongs to a random matrix ensemble is to calculate its energy level statistics and the Kullback-Leibler divergence that captures the level repulsion which is a characteristic feature of a thermalized phase. On the other hand, if the energy level statistics is close to the Poisson distribution, the Hamiltonian will display properties of a nonergodic system. We now show that $H_{\textrm{eff}}$ thermalizes by numerically demonstrating its consistency with random matrix theory.
@@ III.2.1 Random matrix theory treatment of $H_{\mathrm{eff}}$
Figure 5: Plot of noise-averaged level spacing ratio (top) and Kullback-Leibler divergence (bottom) of $H_{\rm eff}$ for $L=10$. Error bars represent standard error of mean for $768$ noise realizations.
@@ III.3 Critical noise strength and its scaling
Having shown that $H_{\mathrm{eff}}$ obeys certain properties of RMT, we now use that to solve for other properties of the noisy algorithm, starting with the critical noise $\delta_{c,\rm gap}$ and its scaling with system size. We start by noting that the algorithm stops giving meaningful answers once the two special states enter into the bulk, i.e., once the gap closes. As evident from Fig 3, this happens near the point where the bulk energies span from $-\pi$ to $\pi$ due to weak dependence of the special state energies on $\delta$. We will study $\delta$-dependence of the special state energies in Sec III.4. We define the noise value at this gap-closing point as the critical noise $\delta_{c,\rm gap}$. As indicated by our numerical findings, the quasi-energy of the noisy Grover operator is well captured by the effective Hamiltonian up to the point of merger of the special states, therefore we will study the eigenvalues of the latter to determine the critical noise and its scaling with system size.
@@
where we used that $\mathcal{E}\left(\epsilon_{j}\epsilon_{k}\right)=\frac{1}{3}\delta_{jk}$. Notice that $E_{0}\sim L$ comes from the overall gate count, $M\sim L^{2}$; by reducing scaling of the gate count, we would modify this system-size dependence. This scaling is corroborated numerically in Fig 6.
@@
It is easy to see that irrespective of the fraction of eigenvalues we consider that span the entire Brillouin zone, the scaling of the critical noise does not change. In Fig.7 we illustrate this scaling by considering $2\sigma$ and $3\sigma$ of the bulk eigenvalues. The parameter $\delta_{c,\mathrm{gap}}$ does not represent a conventional phase transition since it scales as a power law with system size $L$. However, one can define the observable: the late-time-average of occupation in the state $|\bar{x}\rangle$, which vanishes at $\delta=\delta_{c,\mathrm{gap}}$ in a manner characteristic of conventional second-order phase transitions. An alternative perspective is to consider $\delta_{\mathrm{eff}}=L\delta$, which represents the aggregate error rate across all qubits. In terms of this rescaled noise parameter, the transition occurs at $\delta_{\textrm{eff}}=\mathcal{O}(1)$. This type of parameter rescaling is common in both mean-field theory models and nonstandard thermodynamic limits [33].
@@
for $1\leq m,n\leq N$, and where $\mathcal{E}_{\mathrm{GUE}}\left[\cdot\right]$ denotes the average over different realizations of the GUE and $\lambda$ sets the units of energy [34]. As a result of this discrepancy in the scaling of elements, the distribution of the eigenvalues of $H_{\rm eff}$ deviates from Wigner’s semicircle law, instead following a normal distribution
@@
Figure 6: (a) Plot of the sum of squared diagonal and off-diagonal elements of $\tilde{H}_{\rm eff}$ as a function of system size in a log-log plot. Error bars indicate standard error of mean over $512$ noise realizations. The best-fit functions $f(x)=aL^{2}$ (off-diagonal) and $g(L)=bL$ (diagonal) are shown with red and blue dashed lines respectively. (b) $E_{0}$ as a function of system size $L$ calculated using Eq. (28) (red) by averaging over $512$ noise realizations, and is compared with the theoretical prediction given by (31). (c) Distribution of bulk eigenvalues of $H_{\rm eff}$ for $L=10$, shown along with the normal distribution [Eq. (34)] (red) and Wigner’s semicircle (dashed).
Figure 7: Critical noise strength $\delta_{c,\rm gap}$ as a function of system size shown in a log-log plot. The exact critical noise is calculated at the point where the gap between the two special states and the bulk closes and is averaged over $20$ noise realizations. The dashed line on the plot represents the best-fit line for these exact noises. Predicted critical noises are averaged over $512$ noise realizations and are calculated when the two special states merge into $67\%,(-2E_{0},2E_{0})$, and $99.7\%,(-3E_{0},3E_{0})$ of the bulk eigenvalues spanning the Floquet Brillouin zone.
It is worth noting that the exact value of the gap closing between bulk and special states will not happen precisely at $\delta_{c,\rm gap}$ derived above, but rather a slightly smaller value. This is because the edges of the spectrum of $H_{\mathrm{eff}}$ do not scale identically as its variance. For example, conventional local models have an extensive energy range but have energy width scaling as $\sqrt{L}$. However, the energy width $E_{0}\sim L$ is the most important parameter for determining ergodicity, as this is the energy scale on which a finite density of states will come in contact with the special states. The precise scaling of the range of bulk energies, which determines initial gap closing but not ergodicity breaking, will be left for future work.
@@ III.4 Analysis of the special states
Finally, we turn our attention to the two special states. In the absence of noise, the actions of the Grover operator on the states $|\textbf{0}\rangle$ and $|\bar{x}\rangle$ are
$G_{0}|\textbf{0}\rangle$ $=$ $\left(\frac{2}{N}-1\right)|\textbf{0}\rangle+\frac{2\sqrt{N-1}}{N}|\bar{x}\rangle$ (35)
$=$ $-\cos(2\theta)|\textbf{0}\rangle+\sin(2\theta)|\bar{x}\rangle,$ (36)
$G_{0}|\bar{x}\rangle$ $=$ $-\frac{2\sqrt{N-1}}{N}|\textbf{0}\rangle+\left(\frac{2}{N}-1\right)|\bar{x}\rangle$ (37)
$=$ $-\cos(2\theta)|\bar{x}\rangle-\sin(2\theta)|\textbf{0}\rangle.$ (38)
Then we can write the noiseless Grover operator in the $\{|\textbf{0}\rangle,|\bar{x}\rangle\}$ block as
@@
Now, adding back in the systematic noise, we can repeat our analysis using the effective Hamiltonian within first-order perturbation theory and apply it to these special states. Restricting Eq.(25) to the special states manifold by writing it in $|\textbf{0}\rangle$ and $|\bar{x}\rangle$ basis,
$H^{\rm spec}_{\rm eff}\equiv\begin{bmatrix}\langle\textbf{0}|H_{\rm eff}|\textbf{0}\rangle&\langle\textbf{0}|H_{\rm eff}|\bar{x}\rangle\\ \langle\bar{x}|H_{\rm eff}|\textbf{0}\rangle&\langle\bar{x}|H_{\rm eff}|\bar{x}\rangle\end{bmatrix},$ (40)
@@
The eigenvalues of $-2\theta\tau_{y}+\delta H^{\rm spec}_{\rm eff}-\pi$ capture the quasi-energies of two special states, as shown in Fig 8, once again confirming the effectiveness of first-order perturbation theory.
@@
Figure 9: (a) Variance of the coefficients $b_{x},b_{y}$ and $b_{z}$ as a function of system size shown along with their best-fit curves. Data are averaged over $512$ noises. (b) Plot of average energy gap squared $\langle\Delta^{2}\rangle$ obtained using Eq. (43) as a function of system size $L$ for different noise strengths. Data are averaged over $512$ noise realizations.
If $H_{\mathrm{eff}}$ were an $N\times N$ GUE random matrix, then $H^{\rm spec}_{\rm eff}$ would be a $2\times 2$ GUE random matrix, for which $\mathcal{E}(b_{x,y,z}^{2})\sim 1/N$. However, as we have seen that non-GUE scaling of the diagonals exists in $H_{\mathrm{eff}}$, we need to more carefully examine whether GUE scaling holds within the special state manifold. As shown in Fig 9, indeed we do not have exponential scaling of these Pauli coefficients nor the gap. While the off-diagonal terms do decrease exponentially fast $\theta^{2},b^{2}_{x},b^{2}_{y}\sim 1/N$, the diagonal term instead increases as $b^{2}_{z}\sim L$. Putting these scalings into Eq.( 43), we see that $\mathcal{E}\left[\Delta^{2}\right]\sim A/N+B\delta^{2}/N+CL\delta^{2}$ for some order-1 coefficients $A$, $B$, and $C$. Then, for $\delta\gg\delta_{c,\rm comp}\equiv\sqrt{A/(CLN)}\sim 1/\sqrt{L2^{L}}$, the linear $b_{z}$ term dominates over the $\theta^{2}$ term giving a power-law-large gap. While this increases the oscillation frequency within the special states, it comes at the cost of exponentially decreasing the oscillation amplitude. In the limit of $\delta\gg\delta_{c,\rm comp}$, the states $|0\rangle$ and $|\bar{x}\rangle$ become eigenstates of $G^{\rm spec}(\delta)$ and therefore their probabilities remain at
@@
at all times $T$, up to exponentially small oscillations. Since there is no longer an order-1 probability of finding the target state at any times, this implies that the model has lost computing power. Therefore, our special state analysis unfortunately implies that at this much smaller critical noise $\delta_{c,\rm comp}=\mathcal{O}(2^{-L/2})$ we have a “computing transition,” because the algorithm loses computing power for $\delta\apprge\delta_{c,\rm comp}$. By contrast, the gap-closing critical noise $\delta_{c,\rm gap}=\mathcal{O}(L^{-1})$ is much larger, but occurs well after computational power has been lost.
Given that the large coefficient $b_{z}$ in $H^{\rm spec}_{\rm eff}$ implies a special basis for the special state manifold, one might hope to recover an exponentially small gap by randomly rotating the basis, e.g., by choosing a different noise model. However, we now argue that this is not possible for any local or $k$-local noise model. The key point is that the relatively larger magnitude of the $b_{z}$ component compared to $b_{x}$ and $b_{y}$ comes from local distinguishability of the states $|\textbf{0}\rangle$ and $|\bar{x}\rangle$, meaning that generic local (or $k$-local) perturbations in $H_{\mathrm{eff}}$ will have a (random) diagonal component in this basis. By contrast, distinguishing the off-diagonal cat states $|\textbf{0}\rangle+e^{i\varphi}|\bar{x}\rangle$ requires global operators. This implies that $\left|\langle\textbf{0}|H_{\rm eff}|\bar{x}\rangle\right|^{2}\ll\left|\langle\textbf{0}|H_{\rm eff}|\textbf{0}\rangle\right|^{2},\left|\langle\bar{x}|H_{\rm eff}|\bar{x}\rangle\right|^{2}$, which produces $b_{z}^{2}\gg b_{x,y}^{2}$ for generic $k$-local error models.